# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 13.1. Description of $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system of V-trees associated to $G$
Call any subtree of the tree $T^{\bullet}$ whose edge set consists only of V-edges a V-subtree of $T^{\bullet}$. Let $\Pi^{V}$ be the partition of the tree $T$ (in the sense explained in Subsection 2.7) into b-subtrees which correspond to all maximal V-subtrees of $T^{\bullet}$. Take as $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system obtained from $\Theta[{\mathcal{R}}_{G}]$ by consolidation with respect to the partition $\Pi^{V}$, i.e.
@@
We will now describe in more detail the constituent spaces, the peripherals, and the connecting maps of the tree system $\Theta^{V}[{\mathcal{R}}_{G}]$. Suppose first that a b-subtree $S\in\Pi^{V}$ contains a (black) vertex $t$ for which $\Gamma_{t}$ is the extended Whitehead graph, say $\widetilde{W}_{z}$, associated to some virtually free rigid cluster factor $G(z)$ of $G$. It follows from the form of the V-involution in the graphical connecting system ${\mathcal{R}}_{G}$ that the restriction of the system $\Theta[{\mathcal{R}}_{G}]$ to $S$ is then isomorphic (as a tree system) to the V-tree system $\Theta[\widetilde{\mathcal{V}}_{z}]$ of extended Whitehead graphs $\widetilde{W}_{z}$. In particular, for any black vertex $s$ of $S$ the corresponding graph $\Gamma_{s}$ is isomorphic to $\widetilde{W}_{z}$, and the constituent space $K_{S}$ of the consolidated system $\Theta^{V}[{\mathcal{R}}_{G}]$ is homeomorphic (as a space with peripherals) to the V-tree ${\mathcal{X}}_{V}((\widetilde{W}_{z},\tilde{\mathbf{a}}_{z}))$ (where $\tilde{\mathbf{a}}_{z}$ is the V-involution canonically associated to $\widetilde{W}_{z}$).
Consider now the remaining case when a b-subtree $S\in\Pi^{V}$ contains a (black) vertex $t$ for which $\Gamma_{t}$ is some $\theta$-graph $\Xi_{k}$. It follows from the form of the V-involution in the graphical connecting system ${\mathcal{R}}_{G}$ that the restriction of the system $\Theta[{\mathcal{R}}_{G}]$ to $S$ is then isomorphic (as a tree system) to the V-tree system $\Theta[{\mathcal{V}}_{\Xi_{k}}]$ induced by the connecting V-system ${\mathcal{V}}_{\Xi_{k}}$ (i.e. the standard connecting V-system for $\Xi_{k}$, as described in Example 7.11). In particular, for any black vertex $s$ of $S$ the corresponding graph $\Gamma_{s}$ is isomorhic to $\Xi_{k}$, and the constituent space $K_{S}$ of the consolidated system $\Theta^{V}[{\mathcal{R}}_{G}]$ is homeomorphic (as a space with peripherals) to the internally punctured graph $\Xi_{k}$, i.e. to the space with peripherals $(\Xi_{k})^{\circ}_{int}$ (as in Example 7.11).
@@
Recall that if $S\in\Pi^{V}$ is a subtree such that for its black vertices $t$ the graphs $\Gamma_{t}$ are the extended Whitehead graphs, then the constituent space $K_{S}$ of the system $\Theta^{V}[{\mathcal{R}}_{G}]$ has the form of a V-tree of extended Whitehead graphs ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})={\mathcal{X}}_{V}((\widetilde{W}_{z},\tilde{\mathbf{a}}_{z}))$, and that this V-tree can be expressed as the limit of the tree system $\Psi_{S}:=\Psi[\widetilde{\mathcal{V}}_{z}]$ in which the constituent spaces are the core and the arms of ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$, see Lemma 7.12. Fact 7.7 and Lemma 9.6(1) explain then the nature of the core and the arms appearing in this expression. In the remaining case when $S\in\Pi^{V}$ is a subtree with black vertex graphs $\Gamma_{t}$ isomorphic to a $\theta$-graph $\Xi_{k}$ (for some $k$ depending only on $S$), the constituent space $K_{S}$ coincides with the V-tree of $\theta$-graphs ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$, as described in Example 7.11. In particular, in this case the constituent space $K_{S}$ is isomorphic, as a space with peripherals, to the internally punctured $\theta$-graph $(\Xi_{k})^{\circ}_{int}$. Consequently, this space $K_{S}$ can be expressed as the wedge of ($k$ copies of) punctured segments, as described in Example 7.13. For $S$ as in this case, we denote by $\Psi_{S}$ the tree system which corresponds to the just mentioned wedge. Recall that this system $\Psi_{S}$ consists then of precisely one white vertex, $u_{S}$, having degree $k$ in the corresponding underlying tree, and of precisely $k$ black vertices, all adjacent to $u_{S}$, the constituent spaces at which are punctured segments. We call these constituent spaces the wedge arms, and their common peripheral, which is a doubleton, the wedge core.
Since for any $S\in\Pi^{V}$, each peripheral of $K_{S}$ is contained in precisely one arm (or wedge arm), and disjoint with the core (respectively, wedge core), we can naturally view the tree system $\Theta^{V}[{\mathcal{R}}_{G}]$ as obtained by consolidation from another tree system, denoted $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, for which the constituent spaces are the cores, the arms and the wedge arms in all spaces $K_{S}:S\in\Pi^{V}$. The above described systems $\Psi_{S}$ correspond then to the restricted sybsystems of such a new system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ along which one consolidates the latter to obtain back the system $\Theta^{V}[{\mathcal{R}}_{G}]$. We skip the straightforward further details of this construction.
@@
Let $U$ be the set of $G$-orbits in the set of all those edges of the reduced Bowditch JSJ-tree $T^{r}_{G}$ which have a flexible vertex as one of their endpoints. We will use $U$ as a set of types for arms and wedge arms in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, and we will say that a type $\varepsilon\in U$ is orientable (respectively, non-orientable) if and only if the edge stabilizing group $G(e)$ for any $e\in\varepsilon$ is orientable (respectively, non-orientable). Now, let $A$ be any arm in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. Recall that $A$ corresponds to a class (or a sequence of copies) of edges in one of the extended Whitehead graphs $\Gamma_{z}=\widetilde{W}_{z}$, and the set of edges in $\widetilde{W}_{z}$ whose copies are involved in $A$ (which we denote by $E_{A}$) coincides with the preimage $(\lambda_{z}^{+})^{-1}(x)$ for some (precisely one) vertex $x\in N_{z}^{2}$ (see the notation introduced in the fourth paragraph of Section 10). The vertex $x$ as above is adjacent to precisely one vertex $u_{\xi}$ (as described in the same Section 10), and we put $e=[x,u_{\xi}]$. We associate to $A$ (and call the U-type of $A$) the type $\varepsilon=[e]\in U$, i.e. the $G$-orbit of $e$. This orbit clearly belongs to $U$, since $x$ is a flexible vertex of $T^{r}_{G}$ (which follows from the fact that rigid cluster vertices in the tree $T^{r}_{G}$ are isolated, see Fact 3.5). On the other hand, given any wedge arm $A$ in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, we recall that $A$ is composed of punctured edges obtained by puncturing precisely one edge $\sigma$ in precisely one $\theta$-graph $\Gamma_{z}=\Xi_{k(z)}$. (We then put $E_{A}:=\{\sigma\}$.) Since the edges of $\Gamma_{z}$ are in a fixed bijective correspondence with the edges of $T^{r}_{G}$ adjacent to $z$ (via the bijection $\lambda_{z}^{+}$ described in the fourth paragraph of Section 10), we denote by $e_{\sigma}=\lambda_{z}^{+}(\sigma)$ this edge adjacent to $z$ which corresponds to $\sigma$. We associate to $A$ (and call the U-type of A) the type $\varepsilon=[e_{\sigma}]\in U$, i.e. the $G$-orbit of $e_{\sigma}$.
@@
If some U-type $\varepsilon\in U$ is orientable, its orientation is a choice of an orientation for any 2-ended group $G(e)$ with $e\in\varepsilon$ (and such a choice extends, via conjugations in $G$, to the choices of orientations for all groups $G(e)$ with $e\in\varepsilon$). We will now explain how the choice of an orientation of a U-type induces some orientation on any arm or wedge arm equipped with this U-type. If $\varepsilon$ is the U-type of some arm $A$, there is precisely one rigid cluster vertex $z\in Z$ such that $A$ is an arm in a V-tree of extended Whitehead graphs $\Gamma_{z}$. Moreover, there is precisely one $x\in N^{2}_{z}$ such that $A$ is composed of the copies of the edges from $(\lambda_{z}^{+})^{-1}(x)$, and if $u_{\xi}$ denotes the white vertex on the path from $x$ to $z$ then $[u_{\xi},x]\in\varepsilon$. If we fix an orientation of $\varepsilon$, it induces an orientation of the group $G([u_{\xi},x])$, and this induces an order of the doubleton $P=\partial G([u_{\xi},x])=\partial G(u_{\xi})=\partial G([u_{\xi},z])\subset\partial G(z)=C_{z}$, where $C_{z}$ is the core of the V-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$ (for the last equality, compare Lemma 9.6). Note that $P$ coincides with the set of endpoints of precisely one arm $A^{\prime}$ in the V-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$ which is made of punctured copies of the same edges of $\widetilde{W}_{z}$ as the arm $A$. Moreover, the above order of $P$ uniquely determines the order-induced orientations of all edegs of $\widetilde{W}_{z}$ whose copies are involved in $A^{\prime}$ (compare Remark/Definition 7.16, Lemma 8.18 and Remark 9.10(3)). Consequently, the above order-induced orientations of the edges uniquely induce an orientation of the arm $A$, and we will call it the orientation of $A$ induced from the orientation of $\varepsilon$. Obviously, any orientation of $\varepsilon$ induces also uniquely orientations of all peripherals in any arm $A$ of the U-type $\varepsilon$.
@@
Now, suppose that $\varepsilon$ is the U-type of some wedge arm $A$ of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, and that this type is orientable, with one of the orientations fixed. Let $z\in Z$ be this (unique) vertex of type (v2) for which $A$ is an arm in a V-tree of $\theta$-graphs $\Gamma_{z}=\Xi_{k(z)}$. Then there is precisely one $x\in N_{z}$ such that $A$ is composed of the copies of the edge $(\lambda_{z}^{+})^{-1}(x)$. The fixed orientation of the U-type $\varepsilon$ induces then the orientation of the group $G([z,x])$, and this induces the order of the doubleton $\partial G([z,x])=\partial G(z)$. By our convention from Section 10, we have fixed an identification of the set $\partial G(z)$ with the vertex set of the $\theta$-graph $\Gamma_{z}=\Xi_{k(z)}$, so the above orientation of $\varepsilon$ induces an order of the latter vertex set of $\Gamma_{z}$. This in turn induces in the obvious way an orientation of the edge $(\lambda_{z}^{+})^{-1}(x)$ in $\Gamma_{z}$, and since all punctured copies of this edge appearing inside the arm $A$ are oriented consistently, this induces also an orientation of $A$.
@@
We now pass to describing the Y-labels associated to arms and wedge arms. To describe them, we need to look closer at the set $\mathcal{A}$ of E-connections in the graphical connecting system ${\mathcal{R}}_{G}$. Recall that $\mathcal{A}$ is given as the disjoint union of the blocks ${\mathcal{A}}_{y}:y\in Y$, where $Y$ is the set of $G$-orbits in the set of flexible vertices of $T^{r}_{G}$. Each block ${\mathcal{A}}_{y}$ is described in terms of some set $E_{y}$, where the family $E_{y}:y\in Y$ is certain partition of the set of edges of $\Gamma$ (see Section 10). It follows from the descriptions of the blocks ${\mathcal{A}}_{y}$ and the sets $E_{y}$ (in Section 10) that for any arm or wedge arm $A$ the corresponding set $E_{A}$ is contained in precisely one of the sets $E_{y}$; we then say that $y$ is the Y-label associated to the arm $A$. Observe that any two arms (or wedge arms) with the same U-type have also the same Y-label, so that the partition of the set of arms according to the U-types is a refinement of the partition according to the Y-labels. Moreover, it is not hard to see that for any $y\in Y$ the set
@@
Now, consider any label $y\in Y$ such that the flexible factors $G(v):v\in y$ (in the reduced Bowdich’s JSJ splitting of $G$) are orientable. Then to any arm or wedge arm labelled with $y$ we associate the so called #normal-#\##-orientation in the following way. Recall that in our description of the connecting system ${\mathcal{R}}_{G}$ as given in Section 10, we have fixed in advance, in a $G$-invariant way, orientations for all flexible factors $G(v)$ with $v\in y$, and that this choice of orientations led to the assignment of an orientation to any edge $\varepsilon\in E_{y}$, yielding a collection of oriented edges $E_{y}^{\#}$. Note that for any arm (or wedge arm) $A$ labelled with $y$, the copies of edges from $E_{y}$ appearing in this arm, viewed with the orientations as in $E_{y}^{\#}$, are oriented consistently. Consequently, these orientations induce the orientation of the entire arm $A$. Accordingly, we will denote by $A^{\#}$ the arm $A$ equipped with the just described orientation, and we will call this orientation the #normal-#\##-orientation of $A$. We will then call any arm or wedge arm $A$ as above #normal-#\##-oriented. The arms or wedge arms not equipped with ##\##-orientations (i.e. those having the Y-label $y$ such that the groups $G(v):v\in y$ are non-orientable, as flexible factors) will be called #normal-#\##-non-orientable.
The next observation follows directly from the description of the connecting system ${\mathcal{R}}_{G}$, and from the above descriptions of U-types, Y-labels and ##\##-orientations. To justify part (3), one also needs to refer to the density condition (4)(b) in Definition 6.2, as applied to the tree system $\Theta[{\mathcal{R}}_{G}]$ (see also condition (L3) in the construction from the proof of Lemma 6.3). In part (3) we use the following terminological convention: under notation as in the whole Fact 13.1, given a constituent space $K^{*}_{t}$ of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ which is an arm or a wedge arm, and its E-peripheral $P=\varphi^{*}_{[u,t]}(\Sigma^{*}_{u})$, the peripheral adjacent to $P$ is the peripheral $P^{\prime}=\varphi^{*}_{[u,t^{\prime}]}(\Sigma^{*}_{u})$ in the constituent space $K^{*}_{t^{\prime}}$, where $t^{\prime}$ is the unique black vertex of $T_{*}$ adjacent to $u$ and distinct from $t$.
@@
    If $A$ is ##\##-oriented, then $\hat{A}$ is also ##\##-oriented. Moreover, if we induce the ##\##-orientations of $A$ and $\hat{A}$ to the peripherals $\varphi^{*}_{[u,t]}(\Sigma^{*}_{u})\subset A$ and $\varphi^{*}_{[u,\hat{t}]}(\Sigma^{*}_{u})\subset\hat{A}$, then the gluing map $\varphi^{*}_{[u,\hat{t}]}(\varphi^{*}_{[u,t]})^{-1}$ between those peripherals respects the ##\##-orientations, i.e. it reverses the induced orientations of the peripherals.
@@
      the union of all those E-peripherals $P$ of $A$ for which the adjacent to $P$ peripheral $P^{\prime}$ is contained in an arm or a wedge arm of U-type $\varepsilon$, is dense in $A$;

### v2

@@ 13.1. Description of $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system of V-trees associated to $G$
Call any subtree of the tree $T^{\bullet}$ whose edge set consists only of V-edges a V-subtree of $T^{\bullet}$. Let $\Pi^{V}$ be the partition of the tree $T$ (in the sense explained in Section 2.7) into b-subtrees which correspond to all maximal V-subtrees of $T^{\bullet}$. Take as $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system obtained from $\Theta[{\mathcal{R}}_{G}]$ by consolidation with respect to the partition $\Pi^{V}$, i.e.,
@@
We will now describe in more detail the constituent spaces, the peripherals, and the connecting maps of the tree system $\Theta^{V}[{\mathcal{R}}_{G}]$. Suppose first that a b-subtree $S\in\Pi^{V}$ contains a (black) vertex $t$ for which $\Gamma_{t}$ is the extended Whitehead graph, say $\widetilde{W}_{z}$, associated to some virtually free rigid cluster factor $G(z)$ of $G$. It follows from the form of the V-involution in the graphical connecting system ${\mathcal{R}}_{G}$ that the restriction of the system $\Theta[{\mathcal{R}}_{G}]$ to $S$ is then isomorphic (as a tree system) to the V-tree system $\Theta[\widetilde{\mathcal{V}}_{z}]$ of the extended Whitehead graph $\widetilde{W}_{z}$. In particular, for any black vertex $s$ of $S$ the corresponding graph $\Gamma_{s}$ is isomorphic to $\widetilde{W}_{z}$, and the constituent space $K_{S}$ of the consolidated system $\Theta^{V}[{\mathcal{R}}_{G}]$ is homeomorphic (as a space with peripherals) to the V-tree ${\mathcal{X}}_{V}((\widetilde{W}_{z},\tilde{\mathbf{a}}_{z}))$ (where $\tilde{\mathbf{a}}_{z}$ is the V-involution canonically associated to $\widetilde{W}_{z}$).
Consider now the remaining case when a b-subtree $S\in\Pi^{V}$ contains a (black) vertex $t$ for which $\Gamma_{t}$ is some $\theta$-graph $\Xi_{k}$. It follows from the form of the V-involution in the graphical connecting system ${\mathcal{R}}_{G}$ that the restriction of the system $\Theta[{\mathcal{R}}_{G}]$ to $S$ is then isomorphic (as a tree system) to the V-tree system $\Theta[{\mathcal{V}}_{\Xi_{k}}]$ induced by the connecting V-system ${\mathcal{V}}_{\Xi_{k}}$ (i.e. the standard connecting V-system for $\Xi_{k}$, as described in Example 7.11). In particular, for any black vertex $s$ of $S$ the corresponding graph $\Gamma_{s}$ is isomorhic to $\Xi_{k}$, and the constituent space $K_{S}$ of the consolidated system $\Theta^{V}[{\mathcal{R}}_{G}]$ is homeomorphic (as a space with peripherals) to the internally punctured graph $\Xi_{k}$, i.e., to the space with peripherals $(\Xi_{k})^{\circ}_{int}$ (as in Example 7.11).
@@
Recall that if $S\in\Pi^{V}$ is a subtree such that for its black vertices $t$ the graphs $\Gamma_{t}$ are the extended Whitehead graphs, then the constituent space $K_{S}$ of the system $\Theta^{V}[{\mathcal{R}}_{G}]$ has the form of a V-tree of extended Whitehead graphs ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})={\mathcal{X}}_{V}((\widetilde{W}_{z},\tilde{\mathbf{a}}_{z}))$, and that this V-tree can be expressed as the limit of the tree system $\Psi_{S}:=\Psi[\widetilde{\mathcal{V}}_{z}]$ in which the constituent spaces are the core and the arms of ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$, see Lemma 9.6(2) and Lemma 7.12. Fact 7.7 and Lemma 9.6(1) explain then the nature of the core and the arms appearing in this expression. In the remaining case when $S\in\Pi^{V}$ is a subtree with black vertex graphs $\Gamma_{t}$ isomorphic to a $\theta$-graph $\Xi_{k}$ (for some $k$ depending only on $S$), the constituent space $K_{S}$ coincides with the V-tree of $\theta$-graphs ${\mathcal{X}}_{V}({\mathcal{V}}_{\Xi_{k}})$, as described in Example 7.11. In particular, in this case the constituent space $K_{S}$ is isomorphic, as a space with peripherals, to the internally punctured $\theta$-graph $(\Xi_{k})^{\circ}_{int}$. Consequently, this space $K_{S}$ can be expressed as the wedge of ($k$ copies of) punctured segments, $K_{S}=\lim\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$, as described in Example 7.13. For $S$ as in this case, we put $\Psi_{S}=\Psi_{w}[{\mathcal{V}}_{\Xi_{k}}]$, and we recall that this system $\Psi_{S}$ consists of precisely one white vertex, $u_{S}$, having degree $k$ in the corresponding underlying tree, and of precisely $k$ black vertices, all adjacent to $u_{S}$, the constituent spaces at which are punctured segments. We call these constituent spaces the wedge arms, and we call their common peripheral, which is a doubleton, the wedge core.
Since for any $S\in\Pi^{V}$, each peripheral of $K_{S}$ is contained in precisely one arm (or wedge arm), and disjoint with the core (respectively, wedge core), we can naturally view the tree system $\Theta^{V}[{\mathcal{R}}_{G}]$ as obtained by consolidation from another tree system, denoted $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, for which the constituent spaces are the cores, the arms and the wedge arms in all spaces $K_{S}:S\in\Pi^{V}$. The above described systems $\Psi_{S}$ correspond then to the restricted subsystems of such a new system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ along which one consolidates the latter to obtain back the system $\Theta^{V}[{\mathcal{R}}_{G}]$. We skip the straightforward further details of this construction.
@@
Let $U$ be the set of $G$-orbits in the set of all those edges of the reduced Bowditch JSJ-tree $T^{r}_{G}$ which have a flexible vertex as one of their endpoints. We will use $U$ as a set of types for arms and wedge arms in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, and we will say that a type $\varepsilon\in U$ is orientable (respectively, non-orientable) if the edge stabilizing group $G(e)$ for any $e\in\varepsilon$ is orientable (respectively, non-orientable). Now, let $A$ be any arm in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. Recall that $A$ corresponds to a class (or a sequence of copies) of edges in one of the extended Whitehead graphs $\Gamma_{z}=\widetilde{W}_{z}$, and the set of edges in $\widetilde{W}_{z}$ whose copies are involved in $A$ (which we denote by $E_{A}$) coincides with the preimage $(\lambda_{z}^{+})^{-1}(x)$ for some (precisely one) vertex $x\in N_{z}^{2}$ (see the notation introduced in the fourth paragraph of Section 10). The vertex $x$ as above is adjacent to precisely one vertex $u_{\xi}$ (as described in the same Section 10), and we put $e=[x,u_{\xi}]$. We associate to $A$ (and call the U-type of $A$) the type $\varepsilon=[e]\in U$, i.e., the $G$-orbit of $e$. This orbit clearly belongs to $U$, since $x$ is a flexible vertex of $T^{r}_{G}$ (which follows from the fact that rigid cluster vertices in the tree $T^{r}_{G}$ are isolated, see Fact 3.5). Consider now the remaining case when $A$ is a wedge arm in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, and recall that $A$ is composed of punctured edges obtained by puncturing precisely one edge $\sigma$ in precisely one $\theta$-graph $\Gamma_{z}=\Xi_{k(z)}$. (We then put $E_{A}:=\{\sigma\}$.) Since the edges of $\Gamma_{z}$ are in a fixed bijective correspondence with the vertices of $T^{r}_{G}$ adjacent to $z$ (via the bijection $\lambda_{z}^{+}$ described in the fourth paragraph of Section 10), we denote by $e_{\sigma}=[z,\lambda_{z}^{+}(\sigma)]$ this edge joining $z$ to the vertex which corresponds to $\sigma$. We associate to $A$ (and call the U-type of A) the type $\varepsilon=[e_{\sigma}]\in U$, i.e., the $G$-orbit of $e_{\sigma}$.
@@
If some U-type $\varepsilon\in U$ is orientable, its orientation is a choice of an orientation for any 2-ended group $G(e)$ with $e\in\varepsilon$ (and such a choice extends, via conjugations in $G$, to the choices of orientations for all groups $G(e)$ with $e\in\varepsilon$). We will now explain how the choice of an orientation of a U-type induces some orientation on any arm or wedge arm equipped with this U-type. If $\varepsilon$ is the U-type of some arm $A$, there is precisely one rigid cluster vertex $z\in Z$ such that $A$ is an arm in a V-tree of extended Whitehead graphs $\Gamma_{z}$. Moreover, there is precisely one $x\in N^{2}_{z}$ such that $A$ is composed of copies of edges from $(\lambda_{z}^{+})^{-1}(x)$, and if $u_{\xi}$ denotes the white vertex on the path from $x$ to $z$ then $[u_{\xi},x]\in\varepsilon$. If we fix an orientation of $\varepsilon$, it induces an orientation of the group $G([u_{\xi},x])$, and this induces an order of the doubleton $P=\partial G([u_{\xi},x])=\partial G(u_{\xi})=\partial G([u_{\xi},z])\subset\partial G(z)=C_{z}$, where $C_{z}$ is the core of the V-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$ (for the last equality, compare Lemma 9.6). Note that $P$ coincides with the set of endpoints of precisely one arm $A^{\prime}$ in the V-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$ that is made of punctured copies of the same edges of $\widetilde{W}_{z}$ as the arm $A$. Moreover, the above order of $P$ uniquely determines the order-induced orientations of all edges of $\widetilde{W}_{z}$ whose copies are involved in $A^{\prime}$ (compare Remark/Definition 7.16, Lemma 8.18 and Remark 9.10(3)). Consequently, the above order-induced orientations of the edges uniquely induce an orientation of the arm $A$, and we will call it the orientation of $A$ induced from the orientation of $\varepsilon$. Obviously, any orientation of $\varepsilon$ uniquely induces orientations of all peripherals in any arm $A$ of the U-type $\varepsilon$.
@@
Now, suppose that $\varepsilon$ is the U-type of some wedge arm $A$ of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, and that this type is orientable, with one of the orientations fixed. Let $z\in Z$ be the (unique) vertex of type (v2) for which $A$ is an arm in a V-tree of $\theta$-graphs $\Gamma_{z}=\Xi_{k(z)}$. Then there is precisely one $x\in N_{z}$ such that $A$ is composed of copies of the edge $(\lambda_{z}^{+})^{-1}(x)$. The fixed orientation of the U-type $\varepsilon$ induces then an orientation of the group $G([z,x])$, and this induces an order on the doubleton $\partial G([z,x])=\partial G(z)$. By our convention from Section 10, we have fixed an identification of the set $\partial G(z)$ with the vertex set of the $\theta$-graph $\Gamma_{z}=\Xi_{k(z)}$, so the above orientation of $\varepsilon$ induces an order on the vertex set of $\Gamma_{z}$. This in turn induces, in the obvious way, an orientation of the edge $(\lambda_{z}^{+})^{-1}(x)$ in $\Gamma_{z}$, and since all punctured copies of this edge appearing inside the arm $A$ are oriented consistently, this induces also an orientation of $A$.
@@
We now pass to describing the Y-labels associated to arms and wedge arms. To describe them, we need to look closer at the set $\mathcal{A}$ of E-connections in the graphical connecting system ${\mathcal{R}}_{G}$. Recall that $\mathcal{A}$ is given as the disjoint union of the blocks ${\mathcal{A}}_{y}:y\in Y$, where $Y$ is the set of $G$-orbits in the set of flexible vertices of $T^{r}_{G}$. Each block ${\mathcal{A}}_{y}$ is described in terms of some set $E_{y}$, where the family $E_{y}:y\in Y$ is a certain partition of the set of edges of $\Gamma$ (see Section 10). It follows from the descriptions of the blocks ${\mathcal{A}}_{y}$ and the sets $E_{y}$ (in Section 10) that for any arm or wedge arm $A$ the corresponding set $E_{A}$ is contained in precisely one of the sets $E_{y}$; we then say that $y$ is the Y-label associated to the arm $A$. Observe that any two arms (or wedge arms) with the same U-type also have the same Y-label, so that the partition of the set of arms according to the U-types is a refinement of the partition according to the Y-labels. Moreover, it is not hard to see that for any $y\in Y$ the set
@@
Now, consider any label $y\in Y$ such that the flexible factors $G(v):v\in y$ (in the reduced Bowditch JSJ splitting of $G$) are orientable. Then to any arm or wedge arm labelled with $y$ we associate the so called ##\##-orientation in the following way. Recall that in our description of the connecting system ${\mathcal{R}}_{G}$ as given in Section 10, we fixed in advance, in a $G$-invariant way, orientations for all flexible factors $G(v)$ with $v\in y$, and that this choice of orientations led to the assignment of an orientation to any edge $\varepsilon\in E_{y}$, yielding a collection of oriented edges $E_{y}^{\#}$. Note that for any arm (or wedge arm) $A$ labelled with $y$, the copies of edges from $E_{y}$ appearing in this arm, viewed with the orientations as in $E_{y}^{\#}$, are oriented consistently. Consequently, these orientations induce an orientation of the entire arm $A$. Accordingly, we will denote by $A^{\#}$ the arm $A$ equipped with the just described orientation, and we will call this orientation the ##\##-orientation of $A$. We will then call any arm or wedge arm $A$ as above ##\##-oriented. The arms or wedge arms not equipped with ##\##-orientations (i.e. those having the Y-label $y$ such that the groups $G(v):v\in y$ are non-orientable, as flexible factors) will be called ##\##-non-orientable.
The next observation follows directly from the description of the connecting system ${\mathcal{R}}_{G}$, and from the above descriptions of U-types, Y-labels and ##\##-orientations. To justify part (3), one also needs to refer to the density condition (4) in Definition 6.2, as applied to the tree system $\Theta[{\mathcal{R}}_{G}]$ (see also condition (L3) in the construction from the proof of Lemma 6.3). In part (3) we use the following terminological convention: under notation as in the whole of Fact 13.1, given a constituent space $K^{*}_{t}$ of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ which is an arm or a wedge arm, and one of its E-peripherals $P=\varphi^{*}_{[u,t]}(\Sigma^{*}_{u})$, the peripheral adjacent to $P$ is the peripheral $P^{\prime}=\varphi^{*}_{[u,t^{\prime}]}(\Sigma^{*}_{u})$ in the constituent space $K^{*}_{t^{\prime}}$, where $t^{\prime}$ is the unique black vertex of $T_{*}$ adjacent to $u$ and distinct from $t$.
@@
    If $A$ is ##\##-oriented, then $\hat{A}$ is also ##\##-oriented. Moreover, if we induce the ##\##-orientations of $A$ and $\hat{A}$ to the peripherals $\varphi^{*}_{[u,t]}(\Sigma^{*}_{u})\subset A$ and $\varphi^{*}_{[u,\hat{t}]}(\Sigma^{*}_{u})\subset\hat{A}$, then the gluing map $\varphi^{*}_{[u,\hat{t}]}(\varphi^{*}_{[u,t]})^{-1}$ between those peripherals respects the ##\##-orientations, i.e., it reverses the induced orientations of the peripherals.
@@
      the union of all those E-peripherals $P$ of $A$ for which the peripheral $P^{\prime}$ adjacent to $P$ is contained in an arm or a wedge arm of U-type $\varepsilon$, is dense in $A$;